1.6 Write down the formulae for calculating volume for each of the shapes.
Shape Formula
Cylinder

Rectangular solid

Cube

Pyramid


Answer by mananth(16946)


Cylinder Area of base * vertical height

Rectangular solid = L * W * H

Cube = S^3 whwe s is the side of the cube

Pyramid = (1/3) * s^2 *h . s is side of base, h is vertical height